OpenStudy (anonymous):

Which of the following lines has a slope of -1/2? x + 2y = 0 x - 2y = 0 -x + 2y = 0

OpenStudy (jadedry):

There is a basic equation for linear graphs: Y = mx + b where M is the slope, and B is the Y intercept. in this case b is 0, as the line starts at the origin. If you take the first equation , and convert it to mx + b form you will get: \[x+2y = 0\] \[2y = - x\] \[y = -1/2x\] therefore m = -1/2 , therefore the answer is x+2y = 0

OpenStudy (texaschic101):

put your answer choices in y = mx + b form....and then compare the slopes...the slope will be the number in the m position. Example : -x + 2y = 0 2y = x + 0 y = 1/2x + 0/2 the number in the m position is 1/2...so in this one, the slope is 1/2...not the one you want. Continue to check your other answer choices to find the slope and you will eventually find the slope of -1/2.

OpenStudy (anonymous):

Oh, okay. Thank you both very much. :) You've helped a lot.